# The impact of a primordial gravitational wave background on LISA resolvable sources

### v1

@@ II Methodology
  • Bump Low $f$. $A_{\rm Bump}=1.27\times 10^{-7},\,\Delta=0.5,\,f_{*}=2.2\times 10^{-3}$ Hz. We choose a small width $\Delta$ in order for the SGWB to be reasonably peaked. Our choice of $f_{*}$ roughly corresponds to the ISCO frequency of binaries with $M\sim 10^{7}\,M_{\odot}$. As can be seen from Fig. 1, the SGWB peaks around the frequencies where GBs are expected to be detected and has an amplitude comparable to or even larger than their signals. However, since the SGWB signal dies off at higher frequencies, we do not expect it to have a significant effect on SOBHB detections.
  • Bump High $f$. $A_{\rm Bump}=1.27\times 10^{-7},\,\Delta=0.5,\,f_{*}=2.2\times 10^{-2}$ Hz. Same as the previous benchmark, but now the choice of $f_{*}$ roughly corresponds to the ISCO frequency of binaries with $M\sim 10^{6}\,M_{\odot}$. In addition to that, and unlike Bump Low $f$, this signal is expected to impact measurements of SOBHs, leaving almost unaltered detections of GBs.
@@
Figure 3: Relative 68% CL errors on the total mass $M$ [top] and the luminosity distance $d_{L}$ [center] and absolute 68% CL errors on the sky localization $\Omega$ [bottom] from our FIM analysis. We color the region where the percentage error is $>100\%$ with black. We only perform the Fisher analysis for the events that are detected with ${\rm SNR}>10$ in the absence of an SGWB. The errors are shown as function of the fiducial values in the scatter plots, and we also provide histograms to visualize their distribution. Our results are averaged over 15 random realizations of our populations. We plot results for PopIII , Q3delays , and Q3nodelays populations in the first, second, and third rows respectively. Blue, orange, green, and red colors represent our results without SGWB, for PL, Bump High $f$, and Bump Low $f$ respectively.
@@ III.1 Population analysis
To quantify these effects more rigorously, we extend our analysis beyond simple SNR considerations to assess the degradation of parameter estimation due to the SGWB. We employ the Fisher formalism to derive constraints on three key parameters: the total mass $M$, the luminosity distance $d_{L}$, and the sky localization $\Delta\Omega$. Our results are depicted in Fig. 3. These figures compare the (marginalized) error distributions for events with ${\rm SNR}\geq 10$ in the absence and presence of our three benchmark SGWBs. Regions where the forecast relative error exceeds $100\%$ indicate non-detection. For sky localization, we use a non-detectability criterion of $\log_{10}\Delta\Omega\,[{\rm deg}^{2}]\geq 2$. Our results are based on averaging over 15 realizations, to average over Poisson fluctuations and better represent the actual event distribution.
@@
From Fig. 3, the error distribution for the total mass $M$ shifts by an order of magnitude towards larger errors for all three populations. Notably, errors for Q3delays and Q3nodelays populations remain relatively low compared to PopIII , where errors can reach up to $10\%$.
@@
In Fig. 3, the impact of the SGWB is even more pronounced, particularly for the PopIII population. We observe a degradation of the relative error by approximately one order of magnitude. While this poses no significant threat to populations dominated by high-mass events, it severely impacts populations like PopIII where low-mass events are prevalent, pushing many events into non-detection territory.
@@
Turning to the sky localization $\Delta\Omega$, whose forecast uncertainties are shown in he bottom panels of Fig. 3, we find that constraints on this parameter are also significantly degraded by the presence of an SGWB. Even the Q3delays and Q3nodelays populations are affected. For Q3nodelays , nearly half of the events with ${\rm SNR}\geq 10$ are impacted for the Bump Low and High $f$ cases, and even more so for the PL case.
While our Fisher formalism provides a quantitative estimate of these impacts, its assumptions may break down when approaching the detection threshold, as indicated by the red regions in Fig. 3. Despite this limitation, we expect our estimates to provide a reasonable order of magnitude for the errors and an accurate description of their distribution across populations.
@@
Figure 4: Results of simulated Bayesian parameter estimation for the MBHB 1 [left panels] and MBHB 2 [right panels] systems of Table 1. Shown is the posterior distribution on intrinsic parameters, masses, and spins [upper panels] and on the two angles for sky localization, as well as luminosity distance [lower panels]. Contours are shown at 90 % CL. Blue, orange, green, and red colors represent our results without SGWB, for PL, Bump High $f$, and Bump Low $f$ respectively.
@@
$M$ $10^{5}M_{\odot}$ $3\times 10^{6}M_{\odot}$
@@ III.2 In-depth discussion for illustrative injections
To complement the analysis of the previous section, we delve into the detailed parameter degeneracies for a few illustrative events. We employ a direct sampling approach of the likelihood distribution (4) to capture potential non-Gaussian features and multimodalities of the posterior distributions. The parameters for our representative events are listed in Table 1. We select two injections to represent typical events predicted by our populations. While MBHB 1 is more likely found in populations similar to PopIII , it can also appear as an outlier in populations like Q3delays and Q3nodelays , and vice versa for MBHB 2.
@@
Let us start by discussing the intrinsic parameters. We focus on the spin projections along the angular momentum of the binary, $\chi_{-}$, $\chi_{+}$, on the chirp mass $\mathcal{M}_{c}$, and the mass ratio $q$. Visual inspection of the posterior distributions reveals Gaussian-like shapes for most parameters, with slight non-Gaussianities observed in the contours involving the spins. Interestingly, the position of the SGWB bump does not significantly impact the estimation of the spins for both MBHB 1 and 2. On the other hand, compared to Bump High $f$, the Bump Low $f$ scenario severely affects the estimate of the chirp mass for MBHB 1. This can be attributed to the noise level being comparable to the strain of signals like MBHB 1 during their inspiral stage, thereby reducing the ability to constrain the chirp mass of such events. Among the benchmark SGWBs considered, Bump Low $f$ provides the worst-case scenario for constraining the chirp mass of events with relatively small total masses. This situation may deteriorate further for events with masses of the order of $10^{4}M_{\odot}$, where much of the information about the inspiral phase could be obscured by a large SGWB peaked at small frequencies. Except for this parameter, the PL case stands out as the most pessimistic, providing the least favorable constraints, particularly evident for MBHB 2, where constraints worsen by nearly an order of magnitude compared to the no SGWB scenario, diminishing the expected accuracy for such bright signals. Finally, we stress a non-trivial effect occurring in the $q-\chi_{+}$ plane in the MBHB 2 case. We notice that in the High $f$ Bump case, the ellipse is oriented in a different direction with respect to the PL and Low $f$ bump cases. This effect originates from the partial masking of the last part of the GW signal, which effectively modifies the relative weights of the contributions brought by the different frequencies to the total information. Since most information on the spin parameters comes from the very last part of the GW signal, the High $f$ Bump case degrades more substantially the determination of the spins compared to all other parameters.
@@ IV.1 Stellar Origin Black Hole Binaries
As an illustrative case, we consider an event with characteristics similar to GW150914 [63] and discuss the SGWB’s impact on its detectability prospects with LISA. In practice, we choose the masses, $d_{L}$, and spins to match the measured values, pick the initial frequency, say $f_{\rm i}$, such that the signal exits the LISA band at the end after 4 years of data taking, and randomize all the angular variables, i.e., inclination, sky position, and phase. The results of this analysis are shown in Fig. 6. Notice that the peak at low values in the SNR distribution, which appears in all cases, is mostly due to the dependency on $\iota$. Indeed, the SNR is minimal for $\iota=\pi/2$, which corresponds to the peak of the distribution for the number of events.
@@ IV.2 Galactic Binaries
As a last example, we discuss the impact of large SGWBs on GB detections. For this purpose, we use the LISA Data Challenge (LDC) catalog for GBs based on the populations in [71, 72, 73]. As for SOBHBs, we restrict our analysis to SNR considerations. Since the SNR of a specific source can only decrease in the presence of an SGWB, we start by computing the SNR for all the events in the catalog in the case where only the instrumental noise is present (i.e., we neglect the confusion noise) and drop all the events with SNR $<5$, being undetectable even in the most optimistic scenario. We then compute the SNR for the reduced catalog for the “No SGWB” configuration and the three benchmarks discussed throughout this paper. The results of the analysis are shown in the left panel of Fig. 7, where we show variations of the SNR distribution and in its center panel, where we show the percentage loss of GBs with SNR$>5$ as a function of frequency, taking as a reference the No SGWB case. Since most GB signals sit on the low end of the LISA frequency band (see Fig. 1), a bump at low frequency has the greatest impact on the SNR distribution of the sources, whereas a bump at high frequency has the least. Nevertheless, even in the latter case nearly 90% of the GBs are lost at frequencies in the range [3, 15]similar-toabsent315$\sim$[3,\,15]∼ [ 3 , 15 ] mHz. The PL case, while not as severe as the low-frequency bump, would still significantly reduce the number of sources LISA can resolve, decreasing from tens of thousands to hundreds. As for SOBHBs, if the degradation due to a large SGWB is present, it could dramatically threaten the prospects of resolving individual GBs with LISA.
@@ V Discussion and conclusions
  • MBHBs. We conducted a comprehensive study on how the distribution of the SNR and the errors on the binary parameters, estimated via a Fisher analysis, are impacted by an SGWB. We examined three representative types of MBHB populations, each exemplifying different features. We further complemented this with a detailed Bayesian analysis of a pair of representative events, which we identified among the three populations. Our main findings are that a strong SGWB will reduce the number of detectable events, predominantly affecting populations with relatively small masses ($M/M_{\odot}\sim 10^{4}-10^{5}$). The exact number of non-detectable events depends on the spectral shape of the SGWB. We also find that the errors on many binary parameters can increase by up to an order of magnitude. Interestingly, a large SGWB can strongly impact the sky localization, as we have observed through the appearance of bimodal distributions in the event longitude, even for relatively close events that would be well-localized without the SGWB.

### v2

@@ II Methodology
  • Bump Low $f$. $A_{\rm Bump}=1.27\times 10^{-7},\,\Delta=0.5,\,f_{*}=2.2\times 10^{-3}$ Hz. We choose a small width $\Delta$ in order for the SGWB to be reasonably peaked. Our choice of $f_{*}$ roughly corresponds to the ISCO frequency of binaries with $M_{z}\sim 10^{7}\,M_{\odot}$. As can be seen from Fig. 1, the SGWB peaks around the frequencies where GBs are expected to be detected and has an amplitude comparable to or even larger than their signals. However, since the SGWB signal dies off at higher frequencies, we do not expect it to have a significant effect on SOBHB detections.
  • Bump High $f$. $A_{\rm Bump}=1.27\times 10^{-7},\,\Delta=0.5,\,f_{*}=2.2\times 10^{-2}$ Hz. Same as the previous benchmark, but now the choice of $f_{*}$ roughly corresponds to the ISCO frequency of binaries with $M_{z}\sim 10^{6}\,M_{\odot}$. In addition to that, and unlike Bump Low $f$, this signal is expected to impact measurements of SOBHs, leaving almost unaltered detections of GBs.
@@
Figure 3: Relative 68% CL errors on the total mass $M_{z}$ [top] and the luminosity distance $d_{L}$ [center] and absolute 68% CL errors on the sky localization $\Omega$ [bottom] from our FIM analysis. We color the region where the percentage error is $>100\%$ with black. We only perform the Fisher analysis for the events that are detected with ${\rm SNR}>10$ in the absence of an SGWB. The errors are shown as function of the fiducial values in the scatter plots, and we also provide histograms to visualize their distribution. Our results are averaged over 15 random realizations of our populations. We plot results for PopIII, Q3delays, and Q3nodelays populations in the first, second, and third rows respectively. Blue, orange, green, and red colors represent our results without SGWB, for PL, Bump High $f$, and Bump Low $f$ respectively.
@@ III.1 Population analysis
To quantify these effects more rigorously, we extend our analysis beyond simple SNR considerations to assess the degradation of parameter estimation due to the SGWB. We employ the Fisher formalism to derive constraints on three key parameters: the total mass $M_{z}$, the luminosity distance $d_{L}$, and the sky localization $\Delta\Omega$. Our results are depicted in Fig. 3. These figures compare the (marginalized) error distributions for events with ${\rm SNR}\geq 10$ in the absence and presence of our three benchmark SGWBs. Regions where the forecast relative error exceeds $100\%$ indicate non-detection. For sky localization, we use a non-detectability criterion of $\log_{10}\Delta\Omega\,[{\rm deg}^{2}]\geq 2$. Our results are based on averaging over 15 realizations, to average over Poisson fluctuations and better represent the actual event distribution.
@@
From Fig. 3, the error distribution for the total mass $M_{z}$ shifts by an order of magnitude towards larger errors for all three populations. Notably, errors for Q3delays and Q3nodelays populations remain relatively low compared to PopIII, where errors can reach up to $10\%$.
@@
In Fig. 3, the impact of the SGWB is even more pronounced, particularly for the PopIII population. We observe a degradation of the relative error by approximately one order of magnitude. While this poses no significant threat to populations dominated by high-mass events, it severely impacts populations like PopIII where low-mass events are prevalent, pushing many events into non-detection territory. The scatter plot is also useful to understand the trend of the errors on the mass and confirm this argument. For Q3delays and Q3nodelays, more massive events have a worse determination of $M_{z}$ than low-mass ones. Indeed, as can be seen from Fig. 2, larger-mass events in these populations are more distant than lighter ones. As for PopIII, many low-mass events are associated with high redshift, so that many of the scatters, especially in the Bump low $f$ and Power Law cases, are moved to the right, and the peaks of the histogram consequently shift to a larger error.
The error on the distance also varies depending on the population. Faraway sources typically have larger errors. However, while for populations such as Q3delays and Q3nodelays, the error on the distance does not seem to preclude their detectability, many of the high-redshift events in PopIII have a lower mass (see also second column of Fig. 2 in the previous Section), and are therefore much weaker, making the relative error on the distance larger than 100%. The scatter plots reveal a horizontal shift to the right for all populations, almost independent on the distance. This is consistent with the fact that the error on the distance effectively scales with the SNR of the measurement, which is reduced if an SGWB is present (see also discussion below).
@@
Turning to the sky localization $\Delta\Omega$, whose forecast uncertainties are shown in the bottom panels of Fig. 3, we find that constraints on this parameter are also significantly degraded by the presence of an SGWB. Even the Q3delays and Q3nodelays populations are affected. For Q3nodelays, nearly half of the events with ${\rm SNR}\geq 10$ are impacted for the Bump Low and High $f$ cases, and even more so for the PL case.
While our Fisher formalism provides a quantitative estimate of these impacts, its assumptions may break down when approaching the detection threshold (SNR10less-than-or-similar-toabsent10$\lesssim$ 10≲ 10), as indicated by the red points in Fig. 3. For events with SNR10less-than-or-similar-toabsent10$\lesssim$ 10≲ 10 our results should therefore be considered just as rough order of magnitude estimates of the errors, and most conservative errors should be obtained through an MCMC forecast (see next Section). Despite this limitation, we decided to keep these events in Fig. 3, in order to compare the same number of events in the cases with and without SGWB.
@@
Figure 4: Results of simulated Bayesian parameter estimation for the MBHB 1 [left panels] and MBHB 2 [right panels] systems of Table 1. Shown is the posterior distribution on intrinsic parameters, masses, and spins [upper panels] and on the two angles for sky localization, as well as luminosity distance [lower panels]. Contours are shown at 90% CL. Blue, orange, green, and red colors represent our results without SGWB, for PL, Bump High $f$, and Bump Low $f$ respectively. All intrinsic and extrinsic parameters are varied and marginalized over in the analysis. We present them in separate corner plots only for convenience.
@@
$M_{z}$ $10^{5}M_{\odot}$ $3\times 10^{6}M_{\odot}$
@@ III.2 In-depth discussion for illustrative injections
To complement the analysis of the previous section, we delve into the detailed parameter degeneracies, employing a direct sampling approach of the likelihood distribution (4) to capture potential non-Gaussian features and multimodalities of the posterior distributions. For simplicity, we focus on two representative events, whose parameters are listed in Table 1. We select two injections to represent typical events predicted by our populations. While MBHB 1 is more likely found in populations similar to PopIII, it can also appear as an outlier in populations like Q3delays and Q3nodelays, and vice versa for MBHB 2. Furthermore, we choose the parameters of the events in such a way that they develop multimodalities in the distribution of the sky localization upon addition of the SGWB, a feature which is common to many events in our population and that we discuss in depth below.
@@
Let us start by discussing the intrinsic parameters. We focus on the spin projections along the angular momentum of the binary, $\chi_{-}$, $\chi_{+}$, on the redshifted chirp mass $\mathcal{M}_{c}$, and the mass ratio $q$. Visual inspection of the posterior distributions reveals Gaussian-like shapes for most parameters, with slight non-Gaussianities observed in the contours involving the spins. Interestingly, the position of the SGWB bump does not significantly impact the estimation of the spins for both MBHB 1 and 2. On the other hand, compared to Bump High $f$, the Bump Low $f$ scenario severely affects the estimate of the chirp mass for MBHB 1. This can be attributed to the noise level being comparable to the strain of signals like MBHB 1 during their inspiral stage, thereby reducing the ability to constrain the chirp mass of such events. Among the benchmark SGWBs considered, Bump Low $f$ provides the worst-case scenario for constraining the chirp mass of events with relatively small total masses. This situation may deteriorate further for events with masses of the order of $10^{4}M_{\odot}$, where much of the information about the inspiral phase could be obscured by a large SGWB peaked at small frequencies. Except for this parameter, the PL case stands out as the most pessimistic, providing the least favorable constraints, particularly evident for MBHB 2, where constraints worsen by nearly an order of magnitude compared to the no SGWB scenario, diminishing the expected accuracy for such bright signals. Finally, we stress a non-trivial effect occurring in the $q-\chi_{+}$ plane in the MBHB 2 case. We notice that in the High $f$ Bump case, the ellipse is oriented in a different direction with respect to the PL and Low $f$ bump cases. This effect originates from the partial masking of the last part of the GW signal, which effectively modifies the relative weights of the contributions brought by the different frequencies to the total information. Since most information on the spin parameters comes from the very last part of the GW signal, the High $f$ Bump case degrades more substantially the determination of the spins compared to all other parameters.
@@ IV.1 Stellar Origin Black Hole Binaries
As an illustrative case, we consider an event with characteristics similar to GW150914 [63] and discuss the SGWB’s impact on its detectability prospects with LISA. In practice, we choose the masses, $d_{L}$, and spins to match the measured values, pick the initial frequency, say $f_{\rm i}$, such that the signal exits the LISA band after 4 years of data taking, and randomize all the angular variables, i.e., inclination, sky position, and phase. The results of this analysis are shown in Fig. 6. Notice that the peak at low values in the SNR distribution, which appears in all cases, is mostly due to the dependency on $\iota$. Indeed, the SNR is minimal for $\iota=\pi/2$, which corresponds to the peak of the distribution for the number of events.
@@ IV.2 Galactic Binaries
[…]
@@ V Discussion and conclusions
  • MBHBs. We conducted a comprehensive study on how the distribution of the SNR and the errors on the binary parameters, estimated via a Fisher analysis, are impacted by an SGWB. We examined three representative types of MBHB populations, each exemplifying different features. We further complemented this with a detailed Bayesian analysis of a pair of representative events, which we identified among the three populations. Our main findings are that a strong SGWB will reduce the number of detectable events, predominantly affecting populations with relatively small masses ($M_{z}/M_{\odot}\sim 10^{4}-10^{5}$). The exact number of non-detectable events depends on the spectral shape of the SGWB. We also find that the errors on many binary parameters can increase by up to an order of magnitude. Interestingly, a large SGWB can strongly impact the sky localization, as we have observed through the appearance of bimodal distributions in the event longitude, even for relatively close events that would be well-localized without the SGWB.